OpenStudy (anonymous):

315% of what is 82.53?

OpenStudy (anonymous):

Use this: \[\frac{ is }{ of}=\frac{ }{ 100 }\] your percent goes over the 100, but make sure you change it to a decimal. Now plug in what you know and cross multiply

OpenStudy (anonymous):

315/100=3.15 82.53/3.15=X